Records at Hal’s Accounting Services show the following costs for year 1.

Direct materials and supplies $ 40,000
Employee costs 2,900,000
Total overhead 1,300,000


Production was 25,000 billable hours. Fixed overhead was $700,000.

Assuming no change in billable hours in year 2, direct materials and supplies costs are expected to increase by 10 percent. Direct labor costs are expected to increase by 5 percent. Variable overhead per billable hour is expected to remain the same, but fixed overhead is expected to increase by 5 percent.

Required:

a. Year 2 production is expected to be 20,000 billable hours. What are the estimated direct materials, direct labor, variable overhead, and fixed overhead costs for year 2?
b. Determine the total costs per billable hour for year 1 and year 2.
Answer #1

a) calculate following

Direct material (40000/25000*20000)*1.1 35200
Employee costs (2900000/25000*20000)*1.05 2436000
variable overhead (600000/25000*20000) 480000
Fixed overhead (700000*1.05) 735000

b) Total cost per billable hour

Year 1 = 4240000/25000 = 169.60

Year 2 = 3686200/20000 = 184.31